OpenStudy (anonymous):

3(x+2)+4 (x-3)-5 (x-1)

OpenStudy (anonymous):

a ( x + b ) = ax + ab

OpenStudy (anonymous):

- a ( x + b ) = -ax + -ab

OpenStudy (anonymous):

what

OpenStudy (anonymous):

@Jiimy20110 use this formulae to first expand terms then add similar terms:)

OpenStudy (anonymous):

but idk how to do it

OpenStudy (anonymous):

 :)